OpenStudy (anonymous):

suzie's class is made up of 12 boys and 22 girls. Which ration accurately compares the number of girls to the total number of students in the class?

OpenStudy (bahrom7893):

Total number of kids: 12+22 = 34 Girls: 22 Ratio: 22/34 = 11/17

OpenStudy (bahrom7893):

Please click on become a fan if I helped! I really want to get 100 fans!